# PairFlow: Closed-Form Source-Target Coupling for Few-Step Generation in Discrete Flow Models

### v1

@@ 5.2 Molecule Generation
We begin by benchmarking unconditional molecule generation, where models are tasked with generating SMILES strings (weininger1988smiles) that represent molecules. As illustrated in Fig. 2 and Fig. 3, which summarize validity (left), uniqueness (middle), and novelty (right), PairFlow consistently improves upon its base model UDLM (Schiff:2025UDLM), yielding substantial gains in few-step settings. It facilitates 1-step generation on QM9 (ramakrishnan2014quantum), a challenging setting that requires capturing all token-wise dependencies simultaneously. In this case, validity increases from $17.5$ to $223.4$, corresponding to a $12.8\times$ improvement. Similar trends are observed in the 2-step and 4-step settings, with validity improving by $231\%$ and $47.6\%$, respectively. As shown in Fig. 2 (left), this improvement is particularly significant: the 2-step and 4-step validities of PairFlow are comparable to the 4-step and 8-step validities achieved by UDLM (Schiff:2025UDLM). Comparable improvements are also seen on the ZINC-250k (irwin2012zinc) dataset.
Remarkably, PairFlow introduces minimal overhead—less than $2\%$ of the training cost as shown in Tab. 1—and requires no pretrained models, yet achieves performance comparable to, and in some cases surpassing, models distilled from the same base using DCD (Sahoo:2025Duo) and ReDi (Yoo:2025ReDi), both of which rely on pretrained models and finetuning. On both QM9 (ramakrishnan2014quantum) and ZINC-250k (irwin2012zinc)PairFlow consistently outperforms UDLM + ReDi across all few-step settings, achieving substantially higher 2-step validities on QM9 ($232.4$ vs. $416.0$) and ZINC-250k ($75.9$ vs. $146.3$). At the same time, PairFlow matches the performance of UDLM + DCD, with comparable 2-step validities on QM9 ($416.0$ vs. $530.8$). This is particularly notable given that the additional preprocessing cost of PairFlow amounts to only $0.69\%$ on QM9 (ramakrishnan2014quantum) and $6.16\%$ on ZINC-250k (irwin2012zinc), relative to the full cost of DCD (Sahoo:2025Duo). Detailed numerical results with standard deviations are reported in App. D.
Figure 2: Step-wise performance analysis on the QM9 dataset (ramakrishnan2014quantum). Each plot reports the number of valid (left), unique (middle), and novel (right) SMILES strings (weininger1988smiles) out of 1,024 generated samples. Best viewed when zoomed in.
Figure 3: Step-wise performance analysis on the ZINC-250k dataset (irwin2012zinc). Each plot reports the number of valid (left), unique (middle), and novel (right) SMILES strings (weininger1988smiles) out of 1,024 generated samples. Best viewed when zoomed in.
@@ 5.3 Image Generation
Figure 4: Step-wise performance analysis on discretized image datasets. From left to right: FID on MNIST-Binary (lecun2002gradient), FID on CIFAR-10 (krizhevsky2009learning), and Inception Scores (IS) on CIFAR-10. Best viewed when zoomed in.
@@
We further extend our experiments to image domains where each pixel has discretized intensities. As in Sec. 5.2, we evaluate model performance across multiple sampling steps and summarize the results in Fig. 4. Qualitative samples for MNIST-Binary and CIFAR-10 are shown in Fig. 5. Both qualitative and quantitative results show that PairFlow improves the performance of UDLM (Schiff:2025UDLM) and, in few-step settings, achieves performance comparable to DCD (Sahoo:2025Duo).
On MNIST-Binary (Fig. 4, left), PairFlow achieves an FID of $40.59$ in the 1-step setting, equivalently a $68.9\%$ improvement over UDLM (Schiff:2025UDLM). Consistent gains are observed across other few-step settings as well: at 2 steps, FID is reduced by $63.3\%$ ($15.61$ vs. $42.54$), and at 4 steps by $24.4\%$ ($8.51$ vs. $11.25$). On CIFAR-10 (krizhevsky2009learning), where FID (Heusel:2017FID) and IS (Salimans:2016IS) are reported in Fig. 4 (middle and right), PairFlow likewise outperforms the base UDLM, validating the effectiveness of the discovered source–target pairs.
As in molecular generation (Sec. 5.2), PairFlow performs comparable to distillation-based acceleration methods. On MNIST-Binary, it achieves a lower FID than UDLM+DCD in the 1-step setting ($40.59$ vs. $53.84$) and comparable performance at 2 steps ($15.61$ vs. $16.09$). Likewise, PairFlow performs competitively with UDLM+ReDi at 2 steps ($15.61$ vs. $10.36$), while requiring substantially less compute than both. As summarized in Tab. 1, DCD (Sahoo:2025Duo) requires 40 minutes ($T_{\text{DCD}}$) and ReDi (Yoo:2025ReDi) takes 49 minutes, whereas the preprocessing phase of PairFlow completes in just $1.4$ minutes ($T_{\text{{PairFlow}{}}}$), yielding $28.6\times$ and $35\times$ speedups, respectively. On the CIFAR-10 benchmark, both DCD (Sahoo:2025Duo) and ReDi (Yoo:2025ReDi) degrade model performance, as indicated by the higher FID in Fig. 4 (middle) and lower IS in Fig. 4 (right). The results in Tab. 13 suggest that, overall, acceleration methods do not work well on CIFAR-10. We hypothesize that this issue arises from the low performance of the teacher model, which negatively affects the student model when applying acceleration methods. Detailed results are reported in App. D.
@@ 5.4 Distilling Models Trained with Aligned Pairs
While PairFlow alone achieves performance comparable to, or even exceeding, distillation-based techniques (Yoo:2025ReDi; Sahoo:2025Duo), as shown in Sec. 5.2 and Sec. 5.3, we further emphasize that it also serves as a strong initialization for subsequent distillation, yielding even greater performance gains when combined with existing methods. Crucially, this incurs negligible additional cost relative to the overall time required for distillation.
We validate this by distilling PairFlow, trained on QM9 (ramakrishnan2014quantum), ZINC-250k (irwin2012zinc), and MNIST-Binary, using DCD (Sahoo:2025Duo) and ReDi (Yoo:2025ReDi), and comparing their performance against distilled models whose teachers were the base UDLM (Schiff:2025UDLM). As shown in Fig. 6student models distilled from PairFlow, denoted PairFlow+DCD and PairFlow+ReDi, push the frontier of performance previously achieved by distillation-based techniques. For example, on the QM9 dataset (ramakrishnan2014quantum)PairFlow+DCD substantially improves validity over UDLM+DCD ($453.8$ vs. $323$ for 1 step, $685.8$ vs. $530.8$ for 2 steps). A similar trend is observed for PairFlow+ReDi on ZINC-250k (irwin2012zinc), yielding higher scores in both 1-step ($46.3$ vs. $0.7$) and 2-step ($221.5$ vs. $75.9$) generation. Importantly, as summarized in Tab. 1, these gains are achieved at only minimal additional preprocessing cost: $3.15\%$ of the average runtime of distillation on MNIST-Binary, $0.77\%$ on QM9, and $6.42\%$ on ZINC-250k.
@@ Appendix B Experiment details
In Tab. 2, we summarize the hyperparameters used in the experiments presented in Sec. 5, covering both training and finetuning configurations for each dataset. All reported samples were generated using the greedy-tail denoiser described in (Sahoo:2025Duo). We employed an implementation of the closed-form backward velocity that is optimized at the CUDA level.
For the CIFAR-10 dataset, we follow the same setting as baseline (Schiff:2025UDLM). Tab. 4 reports the FID and IS of baseline and PairFlow measured with 1,000 steps and 50K samples, which are consistent with the results reported in Tab. 6 of (Schiff:2025UDLM) and PairFlow outperforms it.
Table 2: Summary of the training settings used in Sec. 5. Specifically, “Sampling Steps” under PairFlow and ReDi (Yoo:2025ReDi) indicate the number of steps taken to generate pairs, “Teacher Update Period” under DCD (Sahoo:2025Duo) denotes the number of fine-tuning iterations between updates, when the teacher model is replaced by the current student model. “# Pairs” under ReDi (Yoo:2025ReDi) denotes the number of pairs for the fine-tuning.
@@
Table 3: FID (Heusel:2017FID) and IS (Salimans:2016IS) of UDLM (Schiff:2025UDLM) and PairFlow on the CIFAR-10 dataset (krizhevsky2009learning).
@@
Table 4: Total Correlation measure with pairs sampled from UDLM (Schiff:2025UDLM) and PairFlow trained on QM9 (ramakrishnan2014quantum).
@@ C.2 Total Correlation analysis of closed-form velocity
As in Sec. 3.2, Yoo:2025ReDi demonstrated that iteratively refining the joint distribution of source-target pairs in discrete flow models improves few-step performance by reducing the total correlation of the model. In this section, we measure the total correlation following their methodology. Specifically, we perform sampling with neural networks, including UDLM (Schiff:2025UDLM) and PairFlow trained on QM9 (ramakrishnan2014quantum), starting from identical initial states $x_{0}$ but with varying random seeds. We randomly select 20,00020{,}000 initial states $x_{0}$, and for each $x_{0}$, we generate 10 samples with a step size of 256. As shown in Tab. 4, PairFlow achieves a lower total correlation, consistent with the improved performance observed in few-step sampling, as discussed above.
@@ Appendix D Detailed Experimental Results
In this section, we provide the detailed experimental results corresponding to those summarized in Sec. 5. For the molecular datasets (QM9 (ramakrishnan2014quantum) and ZINC-250k (irwin2012zinc)), we generate 1,024 samples across varying timesteps and evaluate validity, uniqueness, and novelty. Reported values are averaged over 10 trials, with standard deviations also included. For the image domain, we report FID on MNIST-Binary (lecun2002gradient), and both FID (Heusel:2017FID) and IS (Salimans:2016IS) on CIFAR-10 (krizhevsky2009learning). Detailed experimental settings are provided in Sec. 5.1.
Results on QM9 are presented in Tables 67, and 8, reporting validity, uniqueness, and novelty, respectively. Corresponding results on ZINC-250k are shown in Tables 910, and 11. Finally, results for the image datasets are summarized in Tab. 12 (FID on MNIST-Binary) and Tab. 13 (FID and IS on CIFAR-10). The FID measured on MNIST-Binary (lecun2002gradient), FID and IS measured on CIFAR-10 (krizhevsky2009learning), are summarized in Tab. 12 and Tab. 13, respectively.
Table 6: Validity scores ($\uparrow$) on QM9 (ramakrishnan2014quantum) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 7: Uniqueness scores ($\uparrow$) on QM9 (ramakrishnan2014quantum) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 8: Novelty scores ($\uparrow$) on QM9 (ramakrishnan2014quantum) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 9: Validity scores ($\uparrow$) on ZINC-250k (irwin2012zinc) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 10: Uniqueness scores ($\uparrow$) on ZINC-250k (irwin2012zinc) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 11: Novelty scores ($\uparrow$) on ZINC-250k (irwin2012zinc) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 12: FID ($\downarrow$) on MNIST-Binary (lecun2002gradient) for various methods across different steps. Best values per column are bolded.
@@
Table 13: FID ($\downarrow$) and IS ($\uparrow$) on CIFAR-10 (krizhevsky2009learning) for various methods across different timesteps. Best values per column are bolded.
@@ Appendix E Experiment on Continuous Flow Matching
Table 14: FID of PairFlow on MNIST (lecun2002gradient) with continuous values, measured using FID over 50K samples across various timesteps. Best values are bolded.
@@ E.1 Closed-form Velocity in Continuous Flow Matching
This formulation has already been introduced in previous works (Karras:2022EDM; Bertrand:2025Closed); however, to the best of our knowledge, no prior work has extended this idea to designing couplings for accelerating flow models using the re-flow technique (Liu:2023RF). In the continuous domain, the backward velocity can be obtained directly by flipping the sign of the forward velocity. In contrast, in the discrete domain, the corresponding expression does not converge as $\lim_{t\to 1}$, and thus the backward velocity cannot be employed for sampling starting from data points. Therefore, in this section, we perform experiments using the forward velocity.

### v2

@@ 5.2 Molecule Generation
We begin by benchmarking unconditional molecule generation, where models are tasked with generating SMILES strings (Weininger, 1988) that represent molecules. As illustrated in Fig. 2 and Fig. 3, which summarize validity (left), uniqueness (middle), and novelty (right), PairFlow consistently improves upon its base model UDLM (Schiff et al., 2025), yielding substantial gains in few-step settings. It facilitates 1-step generation on QM9 (Ramakrishnan et al., 2014), a challenging setting that requires capturing all token-wise dependencies simultaneously. In this case, validity increases from $17.5$ to $223.4$, corresponding to a $12.8\times$ improvement. Similar trends are observed in the 2-step and 4-step settings, with validity improving by $231\%$ and $47.6\%$, respectively. As shown in Fig. 2 (left), this improvement is particularly significant: the 2-step and 4-step validities of PairFlow are comparable to the 4-step and 8-step validities achieved by UDLM (Schiff et al., 2025). Comparable improvements are also seen on the ZINC-250k (Irwin et al., 2012) dataset.
Remarkably, PairFlow introduces minimal overhead—less than $2\%$ of the training cost as shown in Tab. 1—and requires no pretrained models, yet achieves performance comparable to, and in some cases surpassing, models distilled from the same base using DCD (Sahoo et al., 2025) and ReDi (Yoo et al., 2025), both of which rely on pretrained models and finetuning. On both QM9 (Ramakrishnan et al., 2014) and ZINC-250k (Irwin et al., 2012)PairFlow consistently outperforms UDLM + ReDi across all few-step settings, achieving substantially higher 2-step validities on QM9 ($232.4$ vs. $416.0$) and ZINC-250k ($75.9$ vs. $146.3$). At the same time, PairFlow matches the performance of UDLM + DCD, with comparable 2-step validities on QM9 ($416.0$ vs. $530.8$). This is particularly notable given that the additional preprocessing cost of PairFlow amounts to only $0.69\%$ on QM9 (Ramakrishnan et al., 2014) and $6.16\%$ on ZINC-250k (Irwin et al., 2012), relative to the full cost of DCD (Sahoo et al., 2025). Detailed numerical results with standard deviations are reported in App. D.
Figure 2: Step-wise performance analysis on the QM9 dataset (Ramakrishnan et al., 2014). Each plot reports the number of valid (left), unique (middle), and novel (right) SMILES strings (Weininger, 1988) out of 1,024 generated samples. Best viewed when zoomed in.
Figure 3: Step-wise performance analysis on the ZINC-250k dataset (Irwin et al., 2012). Each plot reports the number of valid (left), unique (middle), and novel (right) SMILES strings (Weininger, 1988) out of 1,024 generated samples. Best viewed when zoomed in.
@@ 5.3 Image Generation
Figure 4: Step-wise performance analysis on discretized image datasets. From left to right: FID on MNIST-Binary (LeCun et al., 2002), FID on CIFAR-10 (Krizhevsky and others, 2009), and Inception Scores (IS) on CIFAR-10. Best viewed when zoomed in.
@@
We further extend our experiments to image domains where each pixel has discretized intensities. As in Sec. 5.2, we evaluate model performance across multiple sampling steps and summarize the results in Fig. 4. Qualitative samples for MNIST-Binary and CIFAR-10 are shown in Fig. 5. Both qualitative and quantitative results show that PairFlow improves the performance of UDLM (Schiff et al., 2025) and, in few-step settings, achieves performance comparable to DCD (Sahoo et al., 2025).
On MNIST-Binary (Fig. 4, left), PairFlow achieves an FID of $40.59$ in the 1-step setting, equivalently a $68.9\%$ improvement over UDLM (Schiff et al., 2025). Consistent gains are observed across other few-step settings as well: at 2 steps, FID is reduced by $63.3\%$ ($15.61$ vs. $42.54$), and at 4 steps by $24.4\%$ ($8.51$ vs. $11.25$). On CIFAR-10 (Krizhevsky and others, 2009), where FID (Heusel et al., 2017) and IS (Salimans et al., 2016) are reported in Fig. 4 (middle and right), PairFlow likewise outperforms the base UDLM, validating the effectiveness of the discovered source–target pairs.
As in molecular generation (Sec. 5.2), PairFlow performs comparable to distillation-based acceleration methods. On MNIST-Binary, it achieves a lower FID than UDLM+DCD in the 1-step setting ($40.59$ vs. $53.84$) and comparable performance at 2 steps ($15.61$ vs. $16.09$). Likewise, PairFlow performs competitively with UDLM+ReDi at 2 steps ($15.61$ vs. $10.36$), while requiring substantially less compute than both. As summarized in Tab. 1, DCD (Sahoo et al., 2025) requires 40 minutes ($T_{\text{DCD}}$) and ReDi (Yoo et al., 2025) takes 49 minutes, whereas the preprocessing phase of PairFlow completes in just $1.4$ minutes ($T_{\text{{PairFlow}{}}}$), yielding $28.6\times$ and $35\times$ speedups, respectively. On the CIFAR-10 benchmark, both DCD (Sahoo et al., 2025) and ReDi (Yoo et al., 2025) degrade model performance, as indicated by the higher FID in Fig. 4 (middle) and lower IS in Fig. 4 (right). The results in Tab. 13 suggest that, overall, acceleration methods do not work well on CIFAR-10. We hypothesize that this issue arises from the low performance of the teacher model, which negatively affects the student model when applying acceleration methods. Detailed results are reported in App. D.
@@ 5.4 Distilling Models Trained with Aligned Pairs
While PairFlow alone achieves performance comparable to, or even exceeding, distillation-based techniques (Yoo et al., 2025; Sahoo et al., 2025), as shown in Sec. 5.2 and Sec. 5.3, we further emphasize that it also serves as a strong initialization for subsequent distillation, yielding even greater performance gains when combined with existing methods. Crucially, this incurs negligible additional cost relative to the overall time required for distillation.
We validate this by distilling PairFlow, trained on QM9 (Ramakrishnan et al., 2014), ZINC-250k (Irwin et al., 2012), and MNIST-Binary, using DCD (Sahoo et al., 2025) and ReDi (Yoo et al., 2025), and comparing their performance against distilled models whose teachers were the base UDLM (Schiff et al., 2025). As shown in Fig. 6student models distilled from PairFlow, denoted PairFlow+DCD and PairFlow+ReDi, push the frontier of performance previously achieved by distillation-based techniques. For example, on the QM9 dataset (Ramakrishnan et al., 2014)PairFlow+DCD substantially improves validity over UDLM+DCD ($453.8$ vs. $323$ for 1 step, $685.8$ vs. $530.8$ for 2 steps). A similar trend is observed for PairFlow+ReDi on ZINC-250k (Irwin et al., 2012), yielding higher scores in both 1-step ($46.3$ vs. $0.7$) and 2-step ($221.5$ vs. $75.9$) generation. Importantly, as summarized in Tab. 1, these gains are achieved at only minimal additional preprocessing cost: $3.15\%$ of the average runtime of distillation on MNIST-Binary, $0.77\%$ on QM9, and $6.42\%$ on ZINC-250k.
@@ Appendix B Experiment details
In Tab. 2, we summarize the hyperparameters used in the experiments presented in Sec. 5, covering both training and finetuning configurations for each dataset. All reported samples were generated using the greedy-tail denoiser described in (Sahoo et al., 2025). We employed an implementation of the closed-form backward velocity that is optimized at the CUDA level.
For the CIFAR-10 dataset, we follow the same setting as baseline (Schiff et al., 2025). Tab. 4 reports the FID and IS of baseline and PairFlow measured with 1,000 steps and 50K samples, which are consistent with the results reported in Tab. 6 of (Schiff et al., 2025) and PairFlow outperforms it.
Table 2: Summary of the training settings used in Sec. 5. Specifically, “Sampling Steps” under PairFlow and ReDi (Yoo et al., 2025) indicate the number of steps taken to generate pairs, “Teacher Update Period” under DCD (Sahoo et al., 2025) denotes the number of fine-tuning iterations between updates, when the teacher model is replaced by the current student model. “# Pairs” under ReDi (Yoo et al., 2025) denotes the number of pairs for the fine-tuning.
@@
Table 3: FID (Heusel et al., 2017) and IS (Salimans et al., 2016) of UDLM (Schiff et al., 2025) and PairFlow on the CIFAR-10 dataset (Krizhevsky and others, 2009).
@@
Table 4: Total Correlation measure with pairs sampled from UDLM (Schiff et al., 2025) and PairFlow trained on QM9 (Ramakrishnan et al., 2014).
@@ C.2 Total Correlation analysis of closed-form velocity
As in Sec. 3.2, Yoo et al. (2025) demonstrated that iteratively refining the joint distribution of source-target pairs in discrete flow models improves few-step performance by reducing the total correlation of the model. In this section, we measure the total correlation following their methodology. Specifically, we perform sampling with neural networks, including UDLM (Schiff et al., 2025) and PairFlow trained on QM9 (Ramakrishnan et al., 2014), starting from identical initial states $x_{0}$ but with varying random seeds. We randomly select 20,00020{,}000 initial states $x_{0}$, and for each $x_{0}$, we generate 10 samples with a step size of 256. As shown in Tab. 4, PairFlow achieves a lower total correlation, consistent with the improved performance observed in few-step sampling, as discussed above.
@@ Appendix D Detailed Experimental Results
In this section, we provide the detailed experimental results corresponding to those summarized in Sec. 5. For the molecular datasets (QM9 (Ramakrishnan et al., 2014) and ZINC-250k (Irwin et al., 2012)), we generate 1,024 samples across varying timesteps and evaluate validity, uniqueness, and novelty. Reported values are averaged over 10 trials, with standard deviations also included. For the image domain, we report FID on MNIST-Binary (LeCun et al., 2002), and both FID (Heusel et al., 2017) and IS (Salimans et al., 2016) on CIFAR-10 (Krizhevsky and others, 2009). Detailed experimental settings are provided in Sec. 5.1.
Results on QM9 are presented in Tables 67, and 8, reporting validity, uniqueness, and novelty, respectively. Corresponding results on ZINC-250k are shown in Tables 910, and 11. Finally, results for the image datasets are summarized in Tab. 12 (FID on MNIST-Binary) and Tab. 13 (FID and IS on CIFAR-10). The FID measured on MNIST-Binary (LeCun et al., 2002), FID and IS measured on CIFAR-10 (Krizhevsky and others, 2009), are summarized in Tab. 12 and Tab. 13, respectively.
Table 6: Validity scores ($\uparrow$) on QM9 (Ramakrishnan et al., 2014) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 7: Uniqueness scores ($\uparrow$) on QM9 (Ramakrishnan et al., 2014) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 8: Novelty scores ($\uparrow$) on QM9 (Ramakrishnan et al., 2014) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 9: Validity scores ($\uparrow$) on ZINC-250k (Irwin et al., 2012) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 10: Uniqueness scores ($\uparrow$) on ZINC-250k (Irwin et al., 2012) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 11: Novelty scores ($\uparrow$) on ZINC-250k (Irwin et al., 2012) for various methods across different steps. Best values per column are highlighted in bold.
@@
Table 12: FID ($\downarrow$) on MNIST-Binary (LeCun et al., 2002) for various methods across different steps. Best values per column are bolded.
@@
Table 13: FID ($\downarrow$) and IS ($\uparrow$) on CIFAR-10 (Krizhevsky and others, 2009) for various methods across different timesteps. Best values per column are bolded.
@@ Appendix E Experiment on Continuous Flow Matching
Table 14: FID of PairFlow on MNIST (LeCun et al., 2002) with continuous values, measured using FID over 50K samples across various timesteps. Best values are bolded.
@@ E.1 Closed-form Velocity in Continuous Flow Matching
This formulation has already been introduced in previous works (Karras et al., 2022; Bertrand et al., 2025); however, to the best of our knowledge, no prior work has extended this idea to designing couplings for accelerating flow models using the re-flow technique (Liu et al., 2023). In the continuous domain, the backward velocity can be obtained directly by flipping the sign of the forward velocity. In contrast, in the discrete domain, the corresponding expression does not converge as $\lim_{t\to 1}$, and thus the backward velocity cannot be employed for sampling starting from data points. Therefore, in this section, we perform experiments using the forward velocity.